OpenStudy (anonymous):

A consumer charges a $46.92 purchase on a credit card. The card has a daily interest rate of 0.039%. If the consumer pays off the balance at the end of 30 days, how much in interest will be paid for the purchase? a. $0.06 b. $0.18 c. $0.55 d. $5.49

OpenStudy (anonymous):

Assuming that the interest is not compounding, \[\color{green}{\text{Total cost} = \text{Initial purchase price} + \text{Interest}}\] To calculate the interest: \[\color{green}{\text{Interest} = \text{Initial purchase price} \times \frac{0.039}{100} \times 30}\]

OpenStudy (anonymous):

So, \[\color{green}{\text{Total cost} =46.92 + \big(46.92 \times \frac{0.039}{100} \times 30\big)}\]